# Magnetic Field Configurations in Binary Neutron Star Mergers I: Post-merger Remnant and Disk

### v1

@@ III.2.1 Magnetothermal evolution
Figure 5: Two-dimensional mass-weighted histograms describing the magnetic and thermodynamic evolution of the material in simulation POL_UD as a function of time. The top panel shows the mass evolution in the plane ($B$, $s$), and the bottom panel shows ($B$, $\rho$). From left to right, we show different times relative to the merger time. In the top panel, the mass within the remnant (disk) is enclosed within a purple (green) curve. In the bottom panel, the outer boundary of the remnant and disk are marked by horizontal dashed and dot-dashed lines, respectively.
Figure 6: Two-dimensional mass-weighted histograms describing the magnetic and thermodynamic evolution of the material in simulation POL_UD as a function of time. The top panel shows the mass evolution in the plane ($\beta^{-1}$, $s$) and the bottom panel shows ($\beta^{-1}$, $\rho$). From left to right, we show different times relative to the merger time. In the top panel, the mass within the remnant (disk) is enclosed within a purple (green) curve. In the bottom panel, the outer boundary of the remnant and disk are marked by horizontal dashed and dot-dashed lines, respectively.
We investigate the magnetic and thermodynamic evolution of the post-merger remnant and disk by calculating mass-weighted histograms of these properties. Figure 5 displays the mass evolution as a function of magnetic field intensity, entropy per baryon, and rest-mass density for simulation POL_UD. Each column corresponds to a different snapshot in time; from left to right: $t-t_{\rm mer}=-8$, $-2$, $2$, $15$, and $28.8$ ms. The top panel shows how the distribution of mass evolves in the ($B$, $s$) plane, whereas the bottom panel shows the evolution in the ($B$, $\rho$) plane. In the top panels, the mass corresponding to the remnant (disk) material is enclosed with a green (purple) line. The boundaries between these regions in the bottom panel are marked with horizontal dashed and dot-dashed lines of constant $\rho$, respectively. In Fig. 6, we show equivalent histograms but where the magnetic-to-gas pressure ratio $\beta^{-1}$ is used instead of $B$ to characterize the magnetic evolution of the material.
@@
The entropy per baryon also keeps growing at a slow rate during the post-merger phase, though the largest increase occurs immediately after the merger. This large increase is a consequence of the strong shocks produced in the collision, while the growth afterward may indicate slow magnetic reconnection heating. By the end of the simulations, most of the mass in the disk has an entropy per baryon of $s/k_{\rm B}\gtrsim 10$, while the remnant entropy spans a range $s/k_{\rm B}\sim 0.1-5$, with most of the material concentrated at values $s/k_{\rm B}\sim 1-2$.
Figure 7: Two-dimensional mass-weighted histograms of the magnetic and thermodynamic state of the material at the last snapshot for each of the simulations. Each panel shows the mass evolution in the $(B,s)$ plane. The solid lines in the top row delimit the remnant and disk boundaries as defined in Fig. 5; in the bottom row, these boundaries are delimited by the dashed and dot-dashed lines.
Figure 8: Two-dimensional mass histograms describing the magnetic state and thermodynamic evolution of the material at the last snapshot for each of the simulations. Each panel shows the mass evolution in the $(B,\rho)$ plane. The solid lines in the top row delimit the remnant and disk boundaries as defined in Fig. 5; in the bottom row, these boundaries are delimited by the dashed and dot-dashed lines.
To compare the thermodynamic and magnetic field evolution of the different simulations, we show in Figs. 7 and 8 mass histograms for the last snapshot of each. All simulations show similar behavior in their final snapshot, with small deviations, mainly resulting from the different times after the merger that each simulation reached. In particular, in simulations interrupted at earlier times, the magnetic field is still growing in the inner disk and the outer remnant, and the relation between $\rho$ and $B$ is more complicated. This difference should be alleviated at a later time. It is not clear whether the magnetic field will keep increasing toward the saturation value of $B\sim 10^{14}\leavevmode\nobreak\ {\rm G}$ at densities lower than $10^{11}\ {\rm g}\ {\rm cm}^{3}$ over longer timescales.
@@ III.2.2 Time evolution
Figure 10: Average magnetic field as a function of time for runs with and without reflection symmetry. The black curve shows the average magnetic field for all the SR simulations without reflection symmetry, and the gray shaded area delimits one standard deviation to each side of the curve. The green curve shows the average magnetic field for the two HR simulations, whereas the blue and red curves show the average magnetic field for the POL_UU and the BITANT run, respectively.
@@
Figure 13: Average entropy per baryon in the remnant (top row) and the disk (bottom row) as a function of time for various simulations. The black curve is the averaged curve for all the simulations with equal mass NSs and the SFHo EOS, while the gray area encloses one standard deviation to each side of it. The blue and red curves correspond to the STIFF and POL_q=1.2 simulations, respectively, which are the only ones showing significant differences with the other runs.
Figure 14: Baryonic mass in the remnant (top row) and the disk (bottom row) as a function of time for various simulations. The black curve shows the averaged mass for all the SR simulations with equal mass NSs, the SFHo EOS, and a large initial magnetic field, while the gray area encloses one standard deviation to each side of it. The green curve is the averaged mass for the two HR runs. The other curves correspond to the simulations that show a significant deviation from the average curve.
@@
Figure 9 shows the time evolution of the average magnetic field value within the remnant (upper panel) and the disk (lower panel) for all our magnetized simulations. In the remnant, there are large deviations among the different runs. The exponential growth in the remnant due to the KH instability at the interface shear layer is a consistent outcome across the various setups, though there is a large variance in the maximum value of the magnetic field achieved. After this phase, the magnetic field stabilizes for a short period or decreases, to then grow again at a lower rate due to turbulent amplification and finally due to winding. In Paper II, we discuss in more detail the underlying mechanisms of the magnetic field amplification. Here, we note that the BITANT simulation shows a larger and much faster amplification than the other runs; this is reflected in Fig. 10, which shows how this run compares with the average magnetic field evolution of all the other runs, as well as with the POL_UU run, which has the same initial magnetic field but does not assume reflection symmetry across the equatorial plane. This is likely a consequence of the imposed boundary condition, which forces the $B$ field to be predominantly poloidal on the equatorial plane; such a condition, in turn, may help driving a dynamo that converts the generated toroidal field into poloidal. The magnetic field in the remnant for the BITANT run grows even faster than in the HR runs.
@@
The evolution of the magnetic field topology can be addressed by comparing the toroidal and poloidal components of the magnetic field. Figure 11 shows the average of the toroidal-to-poloidal magnetic field energy ratio ($B_{\phi}^{2}/B_{\bf p}^{2}$) for our BNS merger simulations as a function of time for the remnant (top) and the disk (bottom). In the remnant, before the merger, the poloidal component dominates for all simulations except POL_90d, where the initial dipolar field is aligned with the equatorial plane and gives rise to an oscillatory toroidal component when measured from the center of mass of the binary. Immediately after the merger, the toroidal magnetic field energy starts to dominate over the poloidal field energy both in the disk and the remnant and remains dominant until the end of the simulations. All simulations show slow secular growth in the ratio of these two quantities from the merger up to ${\sim}20$ ms post-merger. After this time, the ratio saturates or slightly decreases, indicating a larger relative growth of the poloidal component of the field while the toroidal field reaches a plateau. The remnants show a large variance in the maximum value of this ratio among the simulations, with the purely toroidal runs reaching the maximum value at $\langle B_{\phi}^{2}/B_{\bf p}^{2}\rangle\sim 30$. Simulation POL_q=1.2 consistently shows the lowest ratio for our simulation sample. In the disk, the toroidal-to-poloidal magnetic energy ratio is also $>1$, but its evolution is more chaotic than that in the remnant; the toroidal field dominates by only a factor of ${\sim}2{-}5$ at most.
@@
The evolution of the mean entropy per baryon in the disk and the remnant is shown in Fig. 13. This quantity evolves similarly across the various simulations, provided the EOS and the mass ratio are the same. We have averaged the various curves for all the simulations that share these properties and compared them with the STIFF and the POL_q=1.2 simulation. For the equal-mass SFHo simulations, the entropy per baryon rapidly grows at the merger due to the compression produced in the collision of the NSs and then saturates at $s/k_{\rm B}\sim 12$ in the disk and $s/k_{\rm B}\sim 1.5$ in the remnant. Both POL_q=1.2 and STIFF simulations show a lower entropy per baryon in the disk and higher in the remnant; this is mainly due to the disk being denser and much more massive than in the equal mass SFHo cases, whereas the remnant is less massive.
@@
All the simulations show a rapid increase in the mass of the disk at the merger, followed by a much slower but steady growth until the end of the simulation. Since the disk mass comes from the remnant, the mass of the latter decreases consistently with time. Naturally, POL_q=1.2 shows a larger disk mass, which starts to build up before the merger due to the tidal disruption of the lighter NS. Simulation STIFF also shows a larger disk mass and a lower remnant mass. The stiff EOS simulation produces a larger remnant with less bound material in its outer boundaries, favoring an increase in the disk mass. Interestingly, simulations with a lower or no magnetic field show a larger disk mass. However, there is no clear trend since LOW_B shows a larger mass than NO_B simulation, which may indicate that this difference is due to numerical artifacts and not a real physical process. The measured values for the disk masses are consistent with those found by other authors for similar merger conditions (e.g., Refs. [70, 71]).
@@ III.2.3 Radial and vertical distribution
[…]
@@
Radial (vertical) profiles for simulation POL_UD are shown in the left (right) column of Fig. 15. From top to bottom, we show the rest-mass density $\rho$, angular velocity $\Omega$, absolute value of the magnetic field $|B|$, toroidal-to-poloidal magnetic field ratio $B_{\phi}/B_{\bf p}$, magnetic-to-gas pressure ratio $\beta^{-1}$, and entropy per baryon $s$. In each plot, the darker the line color, the later the time, starting from ${\sim}1\leavevmode\nobreak\ {\rm ms}$ before the merger until the end of the simulation ${\sim}30\leavevmode\nobreak\ {\rm ms}$. The density profile on the equatorial plane initially shows a very oblate post-merger remnant, but it quickly settles to a more stable shape as the outer layers are ejected on the equatorial plane. Similarly, the angular velocity rapidly settles to a differentially rotating profile with $d\Omega/dR>0$ within the core of the remnant, reaching a maximum angular velocity of ${\sim}10\leavevmode\nobreak\ {\rm rad\leavevmode\nobreak\ ms}^{-1}$ at $R\sim 7\leavevmode\nobreak\ {\rm km}$; this corresponds to a rotational period of ${\sim}0.6\leavevmode\nobreak\ {\rm ms}$. At larger radii, the profile slowly transitions to a quasi-Keplerian disk ($\Omega\sim r^{-3/2}$), whose angular velocity profile is overplotted with a black dotted curve.
@@
Figure 16 shows radial and vertical profiles of the above-mentioned quantities at the last time of each of our simulations. As expected, the rest-mass density is not affected by the different magnetic field configurations, but only by the EOS: simulation STIFF produces a less dense and more extended remnant and a denser disk than simulations with the softer SFHo EOS. The EOS also affects the rotational profile for the STIFF simulation, which reaches a peak at a larger radius of ${\sim}9{-}10\leavevmode\nobreak\ {\rm km}$ and a longer maximum rotational period of ${\sim}0.9\leavevmode\nobreak\ {\rm ms}$. Simulation POL_q=1.2 also shows a slightly different rotational profile within the remnant with a lower maximum angular velocity than the equal-mass binaries. In the vertical direction, the trends are similar, with only STIFF producing a different profile from the others due to the more extended remnant.
The magnetic field profiles present larger differences among the simulations, consistent with the findings of the previous section. Most of the simulations show $d\beta^{-1}/dR>0$ as we described above. Simulation POL_90d shows the least magnetized remnant core, but a value of the magnetic field in the disk consistent with the other runs. Since this run also shows the largest $B_{\phi}/B_{\rm p}$ value within the remnant, the low magnetic field corresponds to a much lower poloidal field than the rest of the simulations, which is a consequence of its low initial value. Finally, simulation BITANT shows the largest magnetic field and magnetization in the outer regions of the remnant, despite its last snapshot corresponding to a time ${\sim}15{-}20\leavevmode\nobreak\ {\rm ms}$ earlier than the rest of the simulations.
The vertical profiles of magnetic field quantities also show large differences among the simulations. Simulation BITANT shows the largest magnetic field and magnetization on the equatorial plane due to a large poloidal magnetic field forced by the boundary conditions. It shows a sharp drop at $|z|\sim 10\leavevmode\nobreak\ {\rm km}$, and an increase again above that height to reach the largest value at $|z|\sim 40\leavevmode\nobreak\ {\rm km}$. The other simulations reach a more constant $\beta^{-1}(z)$ profile but with different final average values in the range $\beta^{-1}\sim 10^{-7}-3\times 10^{-5}$. Simulation POL_UU, in a similar way to the behavior in the radial profile, shows the lowest magnetization in the vertical direction, with $\beta^{-1}\sim 3\times 10^{-8}$ for $|z|>10\leavevmode\nobreak\ {\rm km}$.

### v2

@@ III.2.1 Magnetothermal evolution
Figure 6: Two-dimensional mass-weighted histograms depicting the magnetic and thermodynamic evolution of the material in simulation POL-UD as a function of time. The top panel shows the mass evolution in the plane ($B$, $s$), and the bottom panel shows ($B$, $\rho$). From left to right, we show different times relative to the merger time. In the top panel, the mass within the remnant (disk) is enclosed within a purple (green) curve. In the bottom panel, the outer boundaries of the remnant and disk are marked by horizontal dashed and dot-dashed lines, respectively.
Figure 7: Two-dimensional mass-weighted histograms describing the magnetic and thermodynamic evolution of the material in simulation POL-UD as a function of time. The top panel shows the mass evolution in the plane ($\beta^{-1}$, $s$) and the bottom panel shows ($\beta^{-1}$, $\rho$). From left to right, we show different times relative to the merger time. In the top panel, the mass within the remnant (disk) is enclosed within a purple (green) curve. In the bottom panel, the outer boundaries of the remnant and disk are marked by horizontal dashed and dot-dashed lines, respectively.
We investigate the magnetic and thermodynamic evolution of the post-merger remnant and disk by calculating mass-weighted histograms of these properties. Figure 6 displays the mass evolution as a function of magnetic field intensity, entropy per baryon, and rest-mass density for simulation POL-UD. Each column corresponds to a different snapshot in time; from left to right: $t-t_{\rm mer}=-8$, 2-2, 22, $15$, and $28.8$ ms. The top panel shows how the distribution of mass evolves in the ($B$, $s$) plane, whereas the bottom panel shows the evolution in the ($B$, $\rho$) plane. In the top panels, the mass corresponding to the remnant (disk) material is enclosed with a green (purple) line. The boundaries between these regions in the bottom panel are marked with horizontal dashed and dot-dashed lines of constant $\rho$, respectively. In Fig. 7, we show equivalent histograms but where the magnetic-to-gas pressure ratio $\beta^{-1}$ is used instead of $B$ to characterize the magnetic evolution of the material.
@@
The entropy per baryon also keeps growing at a slow rate during the post-merger phase, though the largest increase occurs immediately after the merger. This large increase is a consequence of the strong shocks produced in the collision interface, while the growth afterward may indicate further sustained shocks, for example, between the spiral arms in the disk, or magnetic reconnection. By the end of the simulations, most of the mass in the disk has an entropy per baryon of $s/k_{\rm B}\gtrsim 10$, while the remnant entropy spans a range $s/k_{\rm B}\sim 0.1-5$, with most of the material concentrated at values $s/k_{\rm B}\sim 1-2$.
Figure 8: Two-dimensional mass-weighted histograms of the magnetic and thermodynamic state of the material at $t-t_{\rm mer}=15\leavevmode\nobreak\ {\rm ms}$ for each of the simulations. Each panel shows the mass evolution in the $(B,s)$ plane. The solid lines in the top row delimit the remnant and disk boundaries as defined in Fig. 6.
Figure 9: Two-dimensional mass histograms describing the magnetic state and thermodynamic evolution of the material at $t-t_{\rm mer}=15\leavevmode\nobreak\ {\rm ms}$ for each of the simulations. Each panel shows the mass evolution in the $(B,\rho)$ plane. The horizontal dashed lines delimit the remnant and disk boundaries as defined in Fig. 6.
To compare the thermodynamic and magnetic field evolution of the different simulations, we show in Figs. 8 and 9 mass histograms for the last snapshot of each. All simulations show similar behavior in their final snapshot, with small deviations, mainly resulting from the different times after the merger that each simulation reached. In particular, in simulations interrupted at earlier times, the magnetic field is still growing in the inner disk and the outer remnant, and the relation between $\rho$ and $B$ is more complicated. This difference should be alleviated at a later time. It is not clear whether the magnetic field will keep increasing toward the saturation value of $B\sim 10^{14}\leavevmode\nobreak\ {\rm G}$ at densities lower than $10^{11}\ {\rm g}\ {\rm cm}^{3}$ over longer timescales.
@@ III.2.2 Time evolution
Figure 11: Average magnetic field as a function of time for runs with and without reflection symmetry. The black curve shows the average magnetic field for all the SR simulations without reflection symmetry, and the gray shaded area delimits one standard deviation to each side of the curve. The green curve shows the average magnetic field for the two HR simulations. In contrast, the blue and red curves show the average magnetic field for the POL-UU and the BITANT run, respectively.
@@
Figure 14: Average entropy per baryon in the remnant (top row) and the disk (bottom row) as a function of time for various simulations. The black curve is the averaged curve for all the simulations with equal mass NSs and the SFHo EOS, while the gray area encloses one standard deviation to each side of it. The blue and red curves correspond to the STIFF and POL-Q12 simulations, respectively, which are the only ones showing significant differences with the other runs. The green curve shows the B0 case, which shows no significant difference with the magnetized runs.
Figure 15: Baryonic mass in the remnant (top row) and the disk (bottom row) as a function of time for various simulations. The black curve shows the averaged mass for all the SR simulations with equal mass NSs, the SFHo EOS, and a large initial magnetic field, while the gray area encloses one standard deviation to each side of it. The green curve is the average mass for the two HR runs. The other curves correspond to simulations which a priori may show different disk and remnant masses: STIFF, POL-Q12, B0, LOWB, and BITANT.
@@
Figure 10 shows the time evolution of the average magnetic field value within the remnant (upper panel) and the disk (lower panel) for all our magnetized simulations. In the remnant, there are large deviations among the different runs. The exponential growth in the remnant due to the KH instability at the interface shear layer is a consistent outcome across the various setups, though there is a large variance in the maximum value of the magnetic field achieved. After this phase, the magnetic field stabilizes for a short period or decreases, to then grow again at a lower rate due to turbulent amplification and finally due to winding. In Paper II, we discuss in more detail the underlying mechanisms of the magnetic field amplification. Here, we note that the BITANT simulation shows a larger and much faster amplification than the other runs; this is reflected in Fig. 11, which shows how this run compares with the average magnetic field evolution of all the other runs, as well as with the POL-UU run, which has the same initial magnetic field but does not assume reflection symmetry across the equatorial plane. It seems plausible that the imposed boundary condition–tending to keep the magnetic field $B$ predominantly poloidal on the equatorial plane–could facilitate a feedback process in which toroidal field generated by winding is, at least in part, converted back into poloidal components. In this view, the winding mechanism draws a toroidal component from the poloidal field, while the boundary condition tends to restore a poloidal structure. This is consistent with the amplification during the Kelvin–Helmholtz phase appearing similar in runs with and without bitant symmetry, with differences becoming more noticeable thereafter, when winding effects may start to play a larger role.
@@
The evolution of the magnetic field topology can be addressed by comparing the toroidal and poloidal components of the magnetic field. Figure 12 shows the average of the toroidal-to-poloidal magnetic field energy ratio ($B_{\phi}^{2}/B_{\bf p}^{2}$) for our BNS merger simulations as a function of time for the remnant (top) and the disk (bottom). In the remnant, before the merger, the poloidal component dominates for all simulations except POL-90D, where the initial dipolar field is aligned with the equatorial plane and gives rise to an oscillatory toroidal component when measured from the center of mass of the binary. Immediately after the merger, the toroidal magnetic field energy starts to dominate over the poloidal field energy both in the disk and the remnant and remains dominant until the end of the simulations. All simulations show slow secular growth in the ratio of these two quantities from the merger up to ${\sim}20$ ms post-merger. After this time, the ratio saturates or slightly decreases, indicating a larger relative growth of the poloidal component of the field while the toroidal field reaches a plateau. The remnants show a large variance in the maximum value of this ratio among the simulations, with the purely toroidal runs reaching the maximum value at $\langle B_{\phi}^{2}/B_{\bf p}^{2}\rangle\sim 30$. Simulation POL-Q12 consistently shows the lowest ratio for our simulation sample. In the disk, the toroidal-to-poloidal magnetic energy ratio is also >1>1, but its evolution is more chaotic than that in the remnant; the toroidal field dominates by only a factor of ${\sim}2{-}5$ at most.
@@
The evolution of the average entropy per baryon in the disk and the remnant is shown in Fig. 14. This quantity evolves similarly across the various simulations, provided the EOS and the mass ratio are the same. We have averaged the various curves for all the simulations that share these properties and compared them with the STIFF, the B0, and the POL-Q12 simulation. For the equal-mass merger simulations with SFHo, the entropy per baryon rapidly grows at the merger due to the compression produced in the collision of the NSs and then saturates at $s/k_{\rm B}\sim 12$ in the disk and $s/k_{\rm B}\sim 1.5$ in the remnant. Magnetic fields seem to have a negligible impact on the bulk entropy evolution: the non-magnetized (B0) run closely matches the mass-weighted specific entropy of the magnetic runs.
Compared with the equal-mass SFHo cases, both POL-Q12 and STIFF exhibit systematically lower mass-weighted entropy in the disk ($\Delta s/k_{\rm B}\sim 2$) and marginally higher values in the remnant ($\Delta s/k_{\rm B}\sim 0.2$). The origin of the lower disk entropy differs between these two cases. In POL-Q12, the lower mass star is partially tidally disrupted before coalescence, depositing very-low entropy material ($s/k_{\rm B}<1$) into the nascent disk. In STIFF, the stars (and the post-merger remnant) are less compact, so early post-merger remnant oscillations transfer a larger fraction of relatively low-entropy material ($s/k_{\rm B}\sim 1-2$) into the disk than in the SFHo cases. The rapid addition of this low-entropy mass increases the mass of the disk (see Fig. 15) while reducing its mass-weighted specific entropy. After this transient outflow subsides, continued shock and viscous heating drive the disk entropy back up. The remnant’s mean entropy is less affected because the transferred mass is small compared with the remnant’s much larger mass and therefore has little impact on its mass-weighted average. The evolution and spatial distribution of the entropy per baryon in our runs is broadly consistent with the results of Ref. [78], who additionally showed that neutrino emission/absorption does not significantly affect this quantity.
@@
All the simulations show a rapid increase in the mass of the disk at the merger, followed by a much slower but steady growth until the end of the simulation. Since the disk mass comes from the remnant, the mass of the latter decreases consistently with time. Naturally, POL-Q12 shows a larger disk mass, which starts to build up before the merger due to the tidal disruption of the lighter NS. Simulation STIFF also shows a larger disk mass and a lower remnant mass. The stiff EOS simulation produces a larger remnant with less bound material in its outer boundaries, favoring an increase in the disk mass. Interestingly, simulations with a lower or no magnetic field show a larger disk mass. However, there is no clear trend since LOWB shows a larger mass than B0 simulation, which may indicate that this difference is due to numerical artifacts and not a real physical process. On the other hand, the BITANT run shows a slightly larger disk mass than average; however, the difference is within the uncertainties and not significant. The measured values for the disk masses are consistent with those found by other authors for similar merger conditions (e.g., Refs. [79, 80]).
@@ III.2.3 Radial and vertical distribution
Figure 16: Radial (left column) and vertical (right column) profiles of several quantities for simulation POL-UD. From top to bottom: rest-mass density $\rho$; angular velocity $\Omega$; absolute value of the magnetic field $|B|$; toroidal-to-poloidal magnetic field ratio $B_{\phi}/B_{\rm p}$; magnetic-to-gas pressure ratio ($\beta^{-1}$). Each curve shows radial profiles for a single time of $t-t_{\rm mer}=-1,\leavevmode\nobreak\ 1,\leavevmode\nobreak\ 10,\leavevmode\nobreak\ 20,\leavevmode\nobreak\ {\rm or}\leavevmode\nobreak\ 28.3\leavevmode\nobreak\ {\rm ms}$, where darker colors correspond to later times. The dark dashed curve in the angular velocity radial profile shows a Keplerian profile, $\Omega\sim r^{-3/2}$.
@@
Radial (vertical) profiles for simulation POL-UD are shown in the left (right) column of Fig. 16. From top to bottom, we show the rest-mass density $\rho$, angular velocity $\Omega$, absolute value of the magnetic field $|B|$, toroidal-to-poloidal magnetic field ratio $B_{\phi}/B_{\bf p}$, magnetic-to-gas pressure ratio $\beta^{-1}$, and entropy per baryon $s$. In each plot, the darker the line color, the later the time, starting from ${\sim}1\leavevmode\nobreak\ {\rm ms}$ before the merger until the end of the simulation ${\sim}30\leavevmode\nobreak\ {\rm ms}$. The density profile on the equatorial plane initially shows a very oblate post-merger remnant, but it quickly settles to a more stable shape as the outer layers are ejected on the equatorial plane. Similarly, the angular velocity rapidly settles to a differentially rotating profile with $d\Omega/dR>0$ within the core of the remnant, reaching a maximum angular velocity of ${\sim}10\leavevmode\nobreak\ {\rm rad\leavevmode\nobreak\ ms}^{-1}$ at $R\sim 7\leavevmode\nobreak\ {\rm km}$; this corresponds to a rotational period of ${\sim}0.6\leavevmode\nobreak\ {\rm ms}$. At larger radii, the profile slowly transitions to a quasi-Keplerian disk ($\Omega\sim r^{-3/2}$), whose angular velocity profile is overplotted with a black dotted curve.
@@
Figure 17 shows radial and vertical profiles of the above-mentioned quantities at the last time of each of our simulations. As expected, the rest-mass density is not affected by the different magnetic field configurations, but only by the EOS: simulation STIFF produces a less dense and more extended remnant and a denser disk than simulations with the softer SFHo EOS. The EOS also affects the rotational profile for the STIFF simulation, which reaches a peak at a larger radius of ${\sim}9{-}10\leavevmode\nobreak\ {\rm km}$ and a longer maximum rotational period of ${\sim}0.9\leavevmode\nobreak\ {\rm ms}$. Simulation POL-Q12 also shows a slightly different rotational profile within the remnant with a lower maximum angular velocity than the equal-mass binaries. In the vertical direction, the trends are similar, with only STIFF producing a different profile from the others due to the more extended remnant.
The magnetic field profiles present larger differences among the simulations, consistent with the findings of the previous section. Most of the simulations show $d\beta^{-1}/dR>0$ as we described above. Simulation POL-90D shows the least magnetized remnant core, but a value of the magnetic field in the disk consistent with the other runs. Since this run also shows the largest $B_{\phi}/B_{\rm p}$ value within the remnant, the low magnetic field corresponds to a much lower poloidal field than the rest of the simulations, which is a consequence of its low initial value. Finally, simulation BITANT shows the largest magnetic field and magnetization in the outer regions of the remnant, despite its last snapshot corresponding to a time ${\sim}15{-}20\leavevmode\nobreak\ {\rm ms}$ earlier than the rest of the simulations.
The vertical profiles of magnetic field quantities also show large differences among the simulations. Simulation BITANT shows the largest magnetic field and magnetization on the equatorial plane due to a large poloidal magnetic field forced by the boundary conditions. It shows a sharp drop at $|z|\sim 10\leavevmode\nobreak\ {\rm km}$, and an increase again above that height to reach the largest value at $|z|\sim 40\leavevmode\nobreak\ {\rm km}$. The other simulations reach a more constant $\beta^{-1}(z)$ profile but with different final average values in the range $\beta^{-1}\sim 10^{-7}-3\times 10^{-5}$. Simulation POL-UU, in a similar way to the behavior in the radial profile, shows the lowest magnetization in the vertical direction, with $\beta^{-1}\sim 3\times 10^{-8}$ for $|z|>10\leavevmode\nobreak\ {\rm km}$.